OpenStudy (anonymous):

what is the integral of tan^2 xdx?

OpenStudy (anonymous):

\[\int\limits \tan^2 xdx\]

OpenStudy (amistre64):

since tan^2 = sec^2-1 than might be easier to compute

OpenStudy (anonymous):

ow right! i'm so stupid. i kept wondering how come the integral of tan^x dx is tanx-1. thank you so much! :)

OpenStudy (amistre64):

\[\int sec^2(x)\ dx-\int 1\ dx\]

OpenStudy (amistre64):

youre welcome

OpenStudy (anonymous):

tanx-x **

OpenStudy (amistre64):

yep, +C

OpenStudy (anonymous):

right.:D